# CookAR: Affordance Augmentations in Wearable AR to Support Kitchen Tool Interactions for People with Low Vision

### v1

@@ 7.2. Challenges in AI-powered AR Development
This paper presents several key technical contributions across CV and HCI by constructing the first egocentric kitchen tool affordance dataset, fine-tuning an affordance segmentation model on our dataset, and developing a fully-functional stereo AR system that generates real-time affordance augmentations. Below, we reflect on key technical challenges stemming from both fields.
@@
AI models for real-world use. Although our fine-tuned model outperforms the base model in accurately segmenting affordances of cooking tools, its mAP is still too low to successfully support dynamic activities like cooking in a real-world deployment study. For instance, its AP@75 is 48.6%, meaning in the worst case, about half of all predictions fail to achieve greater than 75% overlap with the ground truth affordance masks. The recognition results could become even worse in real-world use on a wearable AR system due to the users’ constant head motions and unique behaviors. For example, LV users tend to get much closes to view objects than sighted users (Reynolds et al., 2024). This issue highlights the evaluation gap between HCI and AI: a model that demonstrates a reasonable performance under AI evaluation measures (e.g., mAP, AP@50, AP@75) may be appropriate for an in-lab user study but not a more naturalistic environment found in deployment studies. We suggest that when developing AI models, researchers should consider the potential real-world use cases, human needs, and integration to different hardware platforms (e.g., wearable AR) to enable use in practice.
@@
Affordance models and datasets. As opposed to object recognition models and datasets that attract significant attention in AI (Lin et al., 2014; Deng et al., 2009; He et al., 2016; Liu et al., 2021), the research on affordance models and datasets remains nascent. To address this issue, we collected and labeled an affordance image dataset for kitchen tools and fine-tuned an object detection model on the affordance dataset to balance accuracy and speed. However, due to the relatively small scale of the dataset and the model not being designed for affordance, our system encountered affordance-specific issues. For example, our model often struggled to distinguish different handles, as many handles across various cooking tools look similar. While not affecting the mask generation (so that users can still see the correct augmentations), it interfered with the object tracking model supported by the ZED Mini API, leading to flickering and unstable augmentation rendering. To enhance the system’s robustness and affordance recognition capabilities, future AI research should consider the following key areas: (1) developing larger and more diverse affordance datasets. These datasets should capture a wider variety of object interactions and functionalities, allowing the model to learn from a richer set of scenarios; (2) designing models for affordance detection, such as incorporating training objectives that refine object part relationship understanding for better affordance prediction; (3) on the system level, prioritizing improvements to object tracking algorithms to ensure more stable augmentation rendering, especially in dynamic environments where precise object localization is crucial.
@@
System latency. Latency is always a concern for AR systems, especially considering that off-the-shelf AR devices usually do not have sufficient computational power (e.g., GPU) to support real-time CV models. To enable affordance segmentation, our system streams video data between the AR headset and an external server. However, the latency caused by the data transition prevented the overlays from keeping pace with the participants’ head motions, negatively impacting participants’ trust in CookAR’s intelligence and their perceptions of the system’s usability. To address the system latency, efforts are needed in both software development (e.g., real-time AI models) and hardware advancement (e.g., AR devices with powerful GPUs) to increase the usability of AR systems in dynamic real-world activities.
@@ 7.3. Limitations & Future Directions
There are three primary limitations in this work. First, as an initial prototype, we conducted a qualitative user study with a relatively small number of participants to explore usability and solicit reactions to the idea of affordance augmentations in AR. Future work should conduct larger scale studies with more participants and diverse visual conditions. Second, the aforementioned technical challenges and limitations. Third and finally, the current CookAR system provides only one basic affordance augmentation—solid-colored overlays. Building upon the design insights in our study, future work should incorporate more augmentation options (e.g., outlines) and enable more flexible adjustment (e.g., colors, thickness of the outlines), to provide LV users more personalized experience.
@@ 8. Conclusion
In this paper, we introduce CookAR, a wearable AR prototype that can overlay affordance augmentations in real-time to support safe and efficient kitchen tool interactions for people with low vision. To build CookAR, we assembled an egocentric kitchen tool affordance dataset, fine-tuned an RTMDet-Ins-l model on our dataset (i.e., RTMDet-Ins-l-Cook), and leveraged this modified model with a stereo depth camera and an AR headset to generate real-time affordance augmentations in 3D space. In a technical evaluation of RTMDet-Ins-l-Cook, we found that it outperforms the base RTMDet-Ins-l model in accurately segmenting affordance masks of cooking tools. We then evaluated the CookAR prototype in a three-part lab study with LV participants. Findings indicate that participants prefer affordance augmentations over whole object augmentations for kitchen tool interactions. Participants expressed a desire for an augmentation design with a combination of solid-colored and outlined overlays. The grabbable area should be augmented using solid green overlays, while other affordances including hazardous, entry, exit, containment, intersection, and activation areas should be outlined using a contrasting color. Participants also asked for enhanced customization and automatic adapations for affordance augmentation designs. Our work suggests future designs of affordance augmentations to enhance kitchen tool interactions for LV people and advances the state-of-the-art in AI-powered AR experiences as low vision aids.
@@ Acknowledgements.
This work was primarily supported by an NSF Graduate Research Fellowship and NSF CHS #1763199.References
@@
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative Research in Psychology 3, 2 (2006), 77–101. https://doi.org/10.1191/1478088706qp063oa arXiv:https://www.tandfonline.com/doi/pdf/10.1191/1478088706qp063oa
- Braun and Clarke (2019) Virginia Braun and Victoria Clarke. 2019. Reflecting on reflexive thematic analysis. Qualitative Research in Sport, Exercise and Health 11, 4 (2019), 589–597. https://doi.org/10.1080/2159676X.2019.1628806 arXiv:https://doi.org/10.1080/2159676X.2019.1628806

### v2

@@ 7.2. Challenges in AI-powered AR Development
This paper presents several key technical contributions across CV and HCI by constructing the first egocentric kitchen tool affordance dataset, fine-tuning an affordance segmentation model on our dataset, and developing a fully-functional stereo AR system that generates real-time affordance augmentations. However, our study also revealed the impact of technology limitations on user experiences. For example, while finding CookAR promising, participants gave significantly higher Likert scores for the real-world baseline condition (i.e., without CookAR). Below, we reflect on key technical challenges stemming from both fields.
@@
AI models for real-world use. Although our fine-tuned model outperforms the base model, its mAP is still too low to successfully support dynamic activities like cooking in real-world contexts. For instance, its AP@75 is 48.6%, meaning in the worst case, about half of all predictions fail to achieve greater than 75% overlap with the ground truth affordance masks, resulting in misalignment between augmentations and the original objects. The recognition results could become worse during real-world use on AR glasses due to natural human behaviors like users’ constant head motions. For example, LV users tend to get much closer to view objects than sighted users (Reynolds et al., 2024). This issue highlights the evaluation gap between HCI and AI: a model that performs well under AI metrics (e.g., mAP, AP@50, AP@75) may be suitable for in-lab user studies but less so in naturalistic settings. We suggest that when developing AI models, researchers should consider the potential real-world use cases, human needs, and integration to different hardware platforms (e.g., wearable AR) to enable use in practice.
@@
Affordance models and datasets. As opposed to object recognition models and datasets that attract significant attention in AI (Lin et al., 2014; Deng et al., 2009; He et al., 2016; Liu et al., 2021), research on affordance models and datasets remains nascent. To address this issue, we collected and labeled an affordance image dataset for kitchen tools and fine-tuned an object detection model on the affordance dataset to balance accuracy and speed. However, due to the relatively small scale of the dataset and the RTMDet model not being designed for affordance, our system encountered affordance-specific issues. For example, our model often struggled to distinguish different handles, as many handles across various cooking tools look similar. While not affecting the mask generation (allowing users to still see the correct augmentations), it interfered with the object tracking model supported by the ZED Mini API, leading to flickering and unstable augmentation rendering. This is also noticeable in the video supplement.
To enhance system robustness and affordance recognition capabilities, future AI research should consider the following key areas: (1) Developing larger and more diverse affordance datasets. These datasets should capture a wider variety of object interactions and functionalities, allowing the model to learn from a richer set of scenarios; (2) Designing models for affordance detection, such as incorporating training objectives that refine object part relationship understanding for better affordance prediction; and (3) Improving object tracking algorithms to ensure more stable augmentation rendering, especially in dynamic environments where precise object localization is crucial.
@@
System latency. Latency is always a concern for AR systems, especially since off-the-shelf AR devices usually do not have sufficient computational power (e.g., GPU) to support real-time CV. To enable affordance segmentation, our system streams video data between the AR headset and an external server. However, system latency prevented the overlays from keeping pace with the participants’ head motions, negatively impacting their trust in CookAR’s intelligence and perceived system usability. To address system latency, we need advancements in both software (e.g., real-time AI models) and hardware (e.g., AR devices with powerful GPUs), which will also increase the overall usability of AI-powered AR systems in dynamic real-world activities.
@@ 7.3. Limitations & Future Directions
We outline four primary limitations in this work. First, as an initial prototype, we conducted a qualitative user study with a relatively small number of participants to explore usability and solicit reactions to AR-based affordance augmentations. Future work should conduct larger scale studies with more participants and diverse visual conditions. Second, we re-emphasize the aforementioned technical challenges and limitations. Third, the current CookAR system provides only one basic affordance augmentation—solid-colored overlays. Building upon the design insights in our study, future work should incorporate more augmentation options (e.g., outlines) and enable more flexible adjustments (e.g., colors, thickness of the outlines) to provide LV users more personalized experience. Finally, current CookAR system focuses on leveraging CV methods to detect affordances. While some suggested affordances, such as entrance, exit, and activation areas, can be achieved by dataset extension and model fine-tuning, others may not. For example, detecting heated areas may require a thermal sensor and identifying intersection areas can benefit from a LiDAR sensor. Future research should consider additional sensors beyond RGB cameras.
@@ 8. Conclusion
In this paper, we introduce CookAR, a wearable AR system that overlays affordance augmentations in real-time to support safe and efficient kitchen tool interactions for people with low vision. To build CookAR, we assembled an egocentric kitchen tool affordance dataset, fine-tuned an RTMDet-Ins-l model on our dataset (i.e., RTMDet-Ins-l-Cook), and created an AR system with a stereo depth camera to generate real-time affordance augmentations in 3D space. We evaluated CookAR in a three-part lab study with 10 LV participants. Findings indicate participants’ preferences for affordance augmentations over whole object augmentations for tool interactions, as well as revealing seven types of tool affordances and corresponding augmentation designs preferred by LV users. Our work highlights the promise of affordance augmentations in supporting hand-object interactions for LV people and advances state-of-the-art AI-powered AR technology as low vision aids.
@@ Acknowledgements.
This work was supported by an NSF Graduate Research Fellowship and NSF CHS #1763199. We also thank Yang Li, Sieun Kim, and XunMei Liu for their help with our open-sourced content.References
@@
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative Research in Psychology 3, 2 (2006), 77–101. https://doi.org/10.1191/1478088706qp063oa
- Braun and Clarke (2019) Virginia Braun and Victoria Clarke. 2019. Reflecting on reflexive thematic analysis. Qualitative Research in Sport, Exercise and Health 11, 4 (2019), 589–597. https://doi.org/10.1080/2159676X.2019.1628806
@@
Appendix A Part 1 Protocol
In Part 1, participants compared real-world baseline (i.e., using their typical method in daily life), augmentation baseline (i.e., whole object augmentations), and CookAR (i.e., affordance augmentations) in a tool-grabbing task. We asked participants Likert questions and open-ended questions after each condition. After Part 1, we asked additional qualitative questions.
A.1. 7-Point Likert Scale Questionnaire
1.
How effective is the system you just used? Why?
2.
How comfortable are you with seeing these visualizations? Why?
3.
How distracting are these visualizations? Why?
A.2. Post-Condition Open-Ended Questions
1.
Please describe your overall experience with the visual augmentations.
2.
What did you like about this system?
3.
What did you dislike about this system?
4.
(If any) Ask participants about any interactions we observed to be unconventional (e.g., grabbing a knife by the blade)
5.
How can this system be improved?
6.
In what other scenarios do you think this system would be useful?
7.
Do you have any additional comments about this system that we failed to capture?
A.3. Post-Part Open-Ended Questions
1.
Can you compare your experience completing this study task with and without our research prototypes?
2.
Which condition do you prefer to use? Why?
3.
How can this system be improved? Are there any design recommendations you’d like to make?
4.
In which scenarios do you think this system would be useful?
5.
Do you have any additional comments about this study or the systems?
Appendix B Part 2 Protocol
In Part 2, participants completed a full cooking task using CookAR. They then answered open-ended questions about their experience.
1.
Please describe your overall experience cooking with this system.
2.
What did you like about this system?
3.
What did you dislike about this system?
4.
What are some improvements you want to make to the system?
5.
Are there any design recommendations you would like to make?
6.
In which scenarios do you think this system would be useful?
7.
Any additional comments about this study or this system?
Appendix C Part 3 Protocol
In Part 3, participants brainstormed future designs and applications of CookAR. We asked open-ended questions to guide them.
1.
Can you think of your own design that would improve the usability of our prototype?
2.
In what other scenarios besides cooking do you think this kind of a system can be applicable?
3.
Any additional comments about the system and the entire user study?